# RS-ISRefiner: Towards Better Adapting Vision Foundation Models for Interactive Segmentation of Remote Sensing Images

### v1

@@ IV-C Comparative Experiment
4) Comparison of Efficiency. In this subsubsection of comparisons, we have calculated the number of trainable parameters, FLOPs (Floating Point Operations) of the models, as well as the average time consumed per interaction across the five test datasets. It should be noted that when calculated the average consumed time, all models were tested under identical conditions, including input image size, number of test clicks, hardware specifications for test and so on. As shown in Table IV, due to the frozen backbone, the number of trainable parameters in our model is significantly smaller than that of models with ViT-based Vision Foundation Models as backbones which undergo full-parameter training. Consequently, during the mode training process, our model involves minimal gradient computations and thus achieves faster training speed. From a comprehensive perspective of FLOPs and average time consumption, although our model exhibits high FLOPs and increased complexity, as it incorporates an additional adapter on top of the Vision Foundation Model, in terms of average time consumption, our model performs favorably, with a shorter average testing time compared to models such as RITM[2], PiClick[31] and AdaptiveClick[17]. This indicates that our model also holds certain advantages in the generation speed of segmentation results, facilitating efficient segmentation and annotation of remote sensing images.
@@
By comparing the indicators of different fine-tuning methods in the Table V, although our method has slightly higher interactive testing time and computational complexity than other approaches, its trainable parameters (40.11M) are significantly superior to those of Full Fine-Tuning (98.66M). While ensuring a certain degree of parameter update flexibility to adapt to the task, it greatly reduces the parameter scale compared with full-parameter training, achieving a balance between parameter efficiency and task adaptability. This provides a better choice for scenarios that require moderate parameter adjustments to improve performance.
@@
5) Visualization Comparison. As shown in Fig.8, existing methods often suffer from boundary shifts and regional missegmentation in complex remote sensing scenarios. Traditional convolution-based approaches (e.g., RITM[2], GPCIS[49]) tend to have over/under-segmentation with complex backgrounds or blurry boundaries, leading to low IoU scores. ViT-based methods (e.g., SimpleClick[5], PiClick[31], AdaptiveClick[17], MST[32]) perform better on samples with regular boundaries (Row 8) but still missegment targets with complex shapes (Rows 1, 2, 6, 9) or under illumination/shadow interference (Rows 7, 10). While MFP[18] optimizes boundary segmentation by refining probability maps, it struggles with small (Row 4) and slender targets (Row 6). In contrast, our method achieves stable segmentation across all scenarios, with predictions closely aligning with ground truth and significantly fewer errors. Quantitatively, our IoU exceeds 80% in most cases (even over 90% for some complex targets), whereas other methods generally range 60%-80%. This confirms our method has stronger robustness and boundary adherence in interactive remote sensing image segmentation—especially for complex boundaries or severe background interference—effectively reducing errors and ensuring accurate, complete results.
@@ IV-D Ablation Study
To verify the effectiveness and critical of certain components or techniques in the algorithm, we designed four sets of ablation experiments: the role of ViT output features $F_{ViT}$} and high-quality features $F_{HQ}$}, the impact of different values of fusion parameter $\theta$ on the results, the influence of ViT early feature fusion methods on the original mask correction method and the comparison between the original mask modulation method the improved mask modulation method. The experimental conditions remain consistent with those previously used, with the test datasets being iSAID, SandBar and WHUBuilding, and the evaluation metrics being NoC80 and NoC90.
1) The Role of $F_{ViT}$} and $F_{HQ}$}. As can be observed from Table VI, the model achieves the poorest performance when both $F_{ViT}$} and $F_{HQ}$} are absent, indicating that these two features exert a certain influence on the model’s performance. Specifically, the output of Vision Foundation Model possesses global semantic properties. Feeding this output into the model decoder helps supplement the semantic information of remote sensing images. $F_{HQ}$} generated by the high-quality feature extraction module, incorporates convolutional operations, enabling it to capture local detailed features in images. This facilitates more refined learning of target boundaries and shapes by the model. Additionally, it is evident that the degree of impact on the model differs significantly between the absence of $F_{ViT}$} alone and the absence of $F_{HQ}$} alone. The absence of $F_{HQ}$} substantially degrades the model’s performance, whereas the absence of $F_{ViT}$} results in only minor performance loss. This can be attributed to the fact that $F_{ViT}$} output by the ViT, undergoes processing by SimpleFPN to generate $F_{multiscale}$}, with $F_{multiscale}$} retaining most of the semantic information from $F_{ViT}$}. Consequently, the absence of $F_{ViT}$} has a limited impact on model performance. Conversely, $F_{HQ}$} represents an additionally extracted high-quality image feature that holds high value for model’s feature learning. The lack of this feature leads to suboptimal segmentation results for remote sensing targets with complex boundaries.
TABLE VI: Comparison in NoC80 and NoC90 to investigate the effect of $F_{ViT}$} and $F_{HQ}$} on model performance.
@@
TABLE VII: Comparison in NoC80 and NoC90 to analyze the impact of $\theta$ on the model performance.
@@
2) The Impact of $\theta$. Table VII presents the results of ablation experiments on parameter $\theta$. When $\theta$ is set to a fixed value, the overall performance at $\theta=0.5$ outperforms that at $\theta=1$, indicating that $\theta$ can balance the performance contributions of $F_{IFE}$ and $F_{early}$ during cross-attention fusion. Although a fixed $\theta$ is effective for some situations, it struggles to maintain stable performance across different scenarios. In contrast, setting $\theta$ as a learnable parameter yields better results across all datasets and settings. This confirms that the learnable $\theta$ enables the model to adaptively allocate the importance of different feature sources, thereby enhancing overall robustness and generalization capability.
TABLE VIII: Comparison in NoC80 and NoC90 to analyze the impact of $F_{ViT}^{early}$ fusion on the model performance.
@@
Figure 10: Visualization of probability map correction using the original correction algorithm and our proposed algorithm. The left image shows the original input, the middle image presents the correction result of the original algorithm, and the right image illustrates the correction result of our method.
3) The Influence of $F_{ViT}^{early}$ Fusion. Table VIII reports the ablation study on different early ViT feature fusion strategies. Using only a single layer feature $F_{ViT}^{1}$ or $F_{ViT}^{2}$ leads to suboptimal performance, indicating that features from a single level are insufficient to capture the multi-scale information required in complex scenarios. Fixed Fusion brings moderate improvements, but its inability to adaptively weight different feature sources results in limited gains. In contrast, our proposed Gated Fusion consistently achieves the optimal performance across all datasets and settings. These demonstrate that Gated Fusion can effectively learn the relative importance of features from different layers, exploit their complementarity, and thereby improve the adaptability and reliability of the model in handling diverse and challenging remote sensing conditions.
TABLE IX: Comparison of the impact of the original correction algorithm and our proposed algorithm on the model under the NoC80 and NoC90 metrics.
@@
Original 2.24 6.70 4.79 13.21 3.94 9.11
Ours 2.25 6.51 4.54 12.90 3.65 8.42
4) The Contrast of Original and Improved Modulation Method. To validate the effectiveness of the proposed probability map modulation algorithm, we conduct this ablation study. Quantitatively, as shown in Table IX, on the iSAID, SandBar and WHUBuilding datasets with NoC80 and NoC90 as metrics, our method consistently reduces the error compared to the baseline. For instance, on the SandBar datasets, the NoC80 metric decreases from 4.79 to 4.54, and the NoC90 metric decreases from 13.21 to 12.90. From a qualitative perspective, Fig.10 illustrates that the baseline method produces spurious “circular” response regions in the modulated probability maps, whereas our approach generates results that more closely adhere to the true object contours. This demonstrates the ability of the proposed algorithm to effectively suppress wrong regions in probability maps. Overall, the improved modulation algorithm provides higher-quality probability maps to support subsequent interactive steps, thereby facilitating both enhanced segmentation accuracy and improved interaction efficiency.
@@ V Conclusion
In this work, we propose RS-ISRefiner, a novel interactive segmentation framework specifically designed for remote sensing images. By leveraging frozen Vision Foundation Models in combination with the adapter module, our method effectively balances general feature representation with domain-specific adaptation, significantly reducing training costs while improving boundary precision. Furthermore, the integration of a convolution-transformer hybrid attention mechanism enhances the model’s capability to handle large-scale variations and cluttered backgrounds, while the improved probability map modulation algorithm ensures stable propagation of historical interactions for refined segmentation. Extensive experiments on multiple RS datasets, including iSAID, SandBar, NWPU, LoveDA Urban and WHUBuilding, demonstrate that RS-ISRefiner consistently outperforms state-of-the-art approaches in terms of segmentation accuracy, interaction efficiency, and robustness. In particular, our method achieves superior performance under high-precision settings (NoC@85 and NoC@90), validating its effectiveness in practical annotation scenarios.
In future work, we plan to investigate more efficient fine-tuning strategies to further reduce computational complexity, making our framework more suitable for large-scale deployment in operational remote sensing scenarios. We also aim to extend RS-ISRefiner to more challenging settings, such as multimodal remote sensing data (e.g., hyperspectral images), large-scale annotation platforms and real-time human–computer interaction systems.

### v2

@@ IV-C Comparative Experiment
4) Comparison of Efficiency. In this subsubsection of comparisons, we have calculated the number of trainable parameters, FLOPs (Floating Point Operations) of the models, as well as the average time consumed per interaction across the five test datasets. It should be noted that when calculated the average consumed time, all models were tested under identical conditions, including input image size, number of test clicks, hardware specifications for test and so on. As shown in Table III, due to the frozen backbone, the number of trainable parameters in our model is significantly smaller than that of models with ViT-based Vision Foundation Models as backbones which undergo full-parameter training. Consequently, during the mode training process, our model involves minimal gradient computations and thus achieves faster training speed. From a comprehensive perspective of FLOPs and average time consumption, although our model exhibits high FLOPs and increased complexity, as it incorporates an additional adapter on top of the Vision Foundation Model, in terms of average time consumption, our model performs favorably, with a shorter average testing time compared to models such as RITM[2], PiClick[35] and AdaptiveClick[21]. This indicates that our model also holds certain advantages in the generation speed of segmentation results, facilitating efficient segmentation and annotation of remote sensing images.
@@
By comparing the indicators of different fine-tuning methods in the Table IV, although our method has slightly higher interactive testing time and computational complexity than other approaches, its trainable parameters (40.11M) are significantly superior to those of Full Fine-Tuning (98.66M). While ensuring a certain degree of parameter update flexibility to adapt to the task, it greatly reduces the parameter scale compared with full-parameter training, achieving a balance between parameter efficiency and task adaptability. This provides a better choice for scenarios that require moderate parameter adjustments to improve performance.
@@
5) Visualization Comparison. As shown in Fig.8, existing methods often suffer from boundary shifts and regional missegmentation in complex remote sensing scenarios. Traditional convolution-based approaches (e.g., RITM[2], GPCIS[52]) tend to have over/under-segmentation with complex backgrounds or blurry boundaries, leading to low IoU scores. ViT-based methods (e.g., SimpleClick[11], PiClick[35], AdaptiveClick[21], MST[36]) perform better on samples with regular boundaries (Row 8) but still missegment targets with complex shapes (Rows 1, 2, 6, 9) or under illumination/shadow interference (Rows 7, 10). While MFP[22] optimizes boundary segmentation by refining probability maps, it struggles with small (Row 4) and slender targets (Row 6). In contrast, our improved modulation method achieves stable segmentation across all scenarios, with predictions closely aligning with ground truth and significantly fewer errors. Quantitatively, our IoU exceeds 80% in most cases (even over 90% for some complex targets), whereas other methods generally range from 60% to 80%. This confirms our method has stronger robustness and boundary adherence in interactive remote sensing image segmentation-especially for complex boundaries or severe background interference-effectively reducing errors and ensuring accurate, complete results.
@@ IV-D Ablation Study
To verify the effectiveness and critical of certain components or techniques in the algorithm, we designed four sets of ablation experiments: the role of ViT output features $F_{ViT}$} and high-quality features $F_{HQ}$}, the impact of different values of fusion parameter $\theta$ in Fig.4 on the results, the influence of ViT early feature fusion methods on the original mask correction method and the comparison between the original mask modulation method the improved mask modulation method. The experimental conditions remain consistent with those previously used, with the test datasets being iSAID, SandBar and WHUBuilding, and the evaluation metrics being NoC80 and NoC90.
1) The Role of $F_{ViT}$} and $F_{HQ}$}. As can be observed from Table V, the model achieves the poorest performance when both $F_{ViT}$} and $F_{HQ}$} are absent, indicating that these two features exert a certain influence on the model’s performance. Specifically, the output of Vision Foundation Model possesses global semantic properties. Feeding this output into the model decoder helps supplement the semantic information of remote sensing images. $F_{HQ}$} generated by the high-quality feature extraction module, incorporates convolutional operations, enabling it to capture local detailed features in images. This facilitates more refined learning of target boundaries and shapes by the model. Additionally, it is evident that the degree of impact on the model differs significantly between the absence of $F_{ViT}$} alone and the absence of $F_{HQ}$} alone. The absence of $F_{HQ}$} substantially degrades the model’s performance, whereas the absence of $F_{ViT}$} results in only minor performance loss. This can be attributed to the fact that $F_{ViT}$} output by the ViT, undergoes processing by SimpleFPN to generate $F_{multiscale}$}, with $F_{multiscale}$} retaining most of the semantic information from $F_{ViT}$}. Consequently, the absence of $F_{ViT}$} has a limited impact on model performance. Conversely, $F_{HQ}$} represents an additionally extracted high-quality image feature that holds high value for model’s feature learning. The lack of this feature leads to suboptimal segmentation results for remote sensing targets with complex boundaries.
TABLE V: Comparison in NoC80 and NoC90 to investigate the effect of $F_{ViT}$} and $F_{HQ}$} on model performance.
@@
TABLE VI: Comparison in NoC80 and NoC90 to analyze the impact of $\theta$ on the model performance.
@@
2) The Impact of $\theta$. Table VI presents the results of ablation experiments on parameter $\theta$. When $\theta$ is set to a fixed value, the overall performance at $\theta=0.5$ outperforms that at $\theta=1$, indicating that $\theta$ can balance the performance contributions of $F_{IFE}$ and $F_{early}$ during cross-attention fusion. Although a fixed $\theta$ is effective for some situations, it struggles to maintain stable performance across different scenarios. In contrast, setting $\theta$ as a learnable parameter yields better results across all datasets and settings. This confirms that the learnable $\theta$ enables the model to adaptively allocate the importance of different feature sources, thereby enhancing overall robustness and generalization capability.
TABLE VII: Comparison in NoC80 and NoC90 to analyze the impact of $F_{ViT}^{early}$ fusion on the model performance.
@@
Figure 10: Visualization of probability map modulation using the MFP’s algorithm and our proposed algorithm. The left shows the image input, the middle image presents the modulation result of the MFP’s algorithm, and the right image illustrates the modulation result of our method.
3) The Influence of $F_{ViT}^{early}$ Fusion. Table VII reports the ablation study on different early ViT feature fusion strategies. Using only a single layer feature $F_{ViT}^{1}$ or $F_{ViT}^{2}$ leads to suboptimal performance, indicating that features from a single level are insufficient to capture the multi-scale information required in complex scenarios. Fixed Fusion brings moderate improvements, but its inability to adaptively weight different feature sources results in limited gains. In contrast, our proposed Gated Fusion consistently achieves the optimal performance across all datasets and settings. These demonstrate that Gated Fusion can effectively learn the relative importance of features from different layers, exploit their complementarity, and thereby improve the adaptability and reliability of the model in handling diverse and challenging remote sensing conditions.
TABLE VIII: Comparison of the impact of the MFP’s modulation algorithm and our proposed algorithm on the model under the NoC80 and NoC90 metrics.
@@
MFP 2.24 6.70 4.79 13.21 3.94 9.11
VFM-ISRefiner 2.25 6.51 4.54 12.90 3.65 8.42
4) The Contrast of MFP’s and Our Modulation Method. To validate the effectiveness of the proposed probability map modulation algorithm, we conduct this ablation study. Quantitatively, as shown in Table VIII, on the iSAID, SandBar and WHUBuilding datasets with NoC80 and NoC90 as metrics, our method consistently reduces the error compared to the baseline. For instance, on the SandBar datasets, the NoC80 metric decreases from 4.79 to 4.54, and the NoC90 metric decreases from 13.21 to 12.90. From a qualitative perspective, Fig.10 illustrates that the baseline method produces spurious ”circular” response regions in the modulated probability maps, whereas our approach generates results that more closely adhere to the true object contours. This demonstrates the ability of the proposed algorithm to effectively suppress wrong regions in probability maps. Overall, the improved modulation algorithm provides higher-quality probability maps to support subsequent interactive steps, thereby facilitating both enhanced segmentation accuracy and improved interaction efficiency.
@@ V Conclusion
In this work, we propose VFM-ISRefiner, a novel interactive segmentation framework specifically designed for remote sensing images. By incorporating an adapter module alongside frozen Vision Foundation Model, our method effectively balances general feature representation with domain-specific adaptation, significantly improving boundary precision. Furthermore, the integration of a convolution-transformer hybrid attention mechanism enhances the model’s capability to handle large-scale variations and cluttered backgrounds, while the improved probability map modulation algorithm ensures stable propagation of historical interactions for refined segmentation. Extensive experiments on multiple RS datasets, including iSAID, SandBar, NWPU, LoveDA Urban and WHUBuilding, demonstrate that VFM-ISRefiner consistently outperforms state-of-the-art approaches in terms of segmentation accuracy, interaction efficiency, and robustness. In particular, our method achieves superior performance under high-precision settings (NoC85 and NoC90), validating its effectiveness in practical annotation scenarios.
In future work, we plan to investigate more efficient fine-tuning strategies to further reduce computational complexity, making our framework more suitable for large-scale deployment in operational remote sensing scenarios. We also aim to extend VFM-ISRefiner to more challenging settings, such as multimodal remote sensing data (e.g., hyperspectral images), large-scale annotation platforms and real-time human-computer interaction systems.